Write an equation of the line that passes through the given point and is perpendicular to the given line. Your answer should be written in slope-intercept form.
P(−4, 1), y + 2x = −15

Answer by lwsshak3(11628)

Write an equation of the line that passes through the given point and is perpendicular to the given line. Your answer should be written in slope-intercept form.
P(−4, 1), y + 2x = −15
***
Equation of a line: y=mx+b, m=slope, b=y-intercept
y + 2x = −15
y= -2x−15
slope of given line=-2
slope of perpendicular line=1/2 (negative reciprocal)
equation: y=x/2+b
solving for b using coordinates of given point(-4,1) that is on the line
1=-4/2+b
b=3
equation of perpendicular line:
y=x/2+3